Munoz Company is considering adding a new product. The cost accountant has provided the following data: Expected variable cost of manufacturing $ 45 per unit Expected annual fixed manufacturing costs $ 81,000 The administrative vice president has provided the following estimates: Expected sales commission $ 5 per unit Expected annual fixed administrative costs $ 39,000 The manager has decided that any new product must at least break even in the first year. Required Use the equation method and consider each requirement separately. If the sales price is set at $70, how many units must Munoz sell to break even? Munoz estimates that sales will probably be 8,000 units. What sales price per unit will allow the company to break even? Munoz has decided to advertise the product heavily and has set the sales price at $71. If sales are 11,000 units, how much can the company spend on advertising and still break even?
Answer #1

Part 1

Break-even point (in units) = fixed costs / contribution margin per unit

Contribution margin per unit = selling price per unit – variable cost per unit

Variable costs = variable cost of manufacturing+ sales commission = 45+5 = 50

Contribution margin per unit = 70-50 =$20

Fixed costs = annual fixed manufacturing costs+ annual fixed administrative costs = 81000+39000 = $120000

Break-even point (in units) = 120000/20 = 6000 units

Part 2

At breakeven point, the profit is zero

Profit = SP(Q)-VC(Q)-FC

0 = SP(8000) – 50*8000 – 120000

0 = 8000SP – 400000 – 120000

520000 = 8000SP

SP = 520000/8000

SP = $65

Part 3

Profit = SP(Q)-VC(Q)-FC

0 = (71*11000) – (50*11000) – FC

0 = 781000 – 550000 - FC

FC = 231000

FC = annual fixed manufacturing costs+ annual fixed administrative costs + advertising costs

231000 = 81000+39000+ advertising costs

Advertising costs = 231000-120000 = $111000